what temperature gradient must exist in an aluminum rod for it to transmit 8.0 cal per second per cm
1
Expert's answer
2021-07-07T12:50:55-0400

Question

What temperature gradient must exist in aluminium rod for it to transmit 8.0 cal per second per cm.

Solution;

P="\frac {k×∆T×A}{L}"

p is power

k= thermal conductivity of Aluminium

Take k=205W/mK

∆T is temperature gradient.

L is the length.

Given:

"\frac P A" =8.0cal/scm2=33.4864×104W/sm2

"\frac {∆T}{L}=\frac PA\times \frac 1 k"

"\frac {∆T}{L}=\frac {33.4944×10^4}{205}"=1633.87K/m

Answer

∆T/L=1633.87K/m or 1360.87°C/m